# Self-Alignment: Improving Alignment of Cultural Values in LLMs via In-Context Learning

### v1

@@ Conclusion
In this paper, we presented self-alignment, a simple, yet novel, method to improve the alignment of LLMs at inference time with respect to the cultural values that they encode. Self-alignment exploits the ICL abilities of a LLM to adjust the model’s responses such that they better align to the cultural values of a country. We found that this method proves effective across 5 LLMs and across a variety of languages, albeit to a different extent. Moreover, we found evidence that the LLMs can indeed pick up useful cues from the demonstration examples to induce the correct answers. We envision that the ideas behind self-alignment could prove useful for LLMs in practice. In particular, the WVS survey value responses could be replaced by real user responses and automatically be prepended as demonstration examples to each prompt through the system message (Lee et al. 2024). However, future work should first study how this method will affect model responses in more realistic chat scenarios (i.e., when not explicitly prompting the LLM for cultural values).
Limitations
@@
While this method shows promising results, this work is exploratory in that we were only able to evaluate the method in a somewhat artificial, controlled setting, where all samples were presented in the same format and were originally carefully curated by social scientists. In practice, it is unlikely that users will explicitly ask the model to answer questions which directly probe cultural values. As such, this method would only be practically useful if providing these types of demonstration examples will also prompt the model to adjust its responses when talking about cultural values implicitly. And while we show that ICL-based self-alignment is a promising approach in principle, future work should further investigate in what ways this method affects cultural alignment in more realistic chat scenarios.
@@
Moreover, as mentioned in the related work, it is not entirely clear what we should be aligning LLMs to. For this study, we always attempt to align the entire distribution of model responses to the majority answer from the WVS survey results. However, it might also be sensible to aim for a soft alignment to better reflect the human data. For instance, we could allow the LLMs to be less certain about answers to questions for which human responses were more divided.
Appendices
Appendix A A Ethical considerations
@@
Appendix B B Language to country mapping
@@
Appendix C C Model details
We use 5 LLMs of varying sizes of which all are open-source except for Gemini. For our open source model, we rely on the HuggingFace implementation (Wolf et al. 2019), and for Gemini we use Google’s paid API service.
@@ Appendix C C Model details
Similar to Llama3-8B, Mistral AI 7B (Jiang et al. 2023) is English-centric and mostly pretrained on Latin-script languages. For our experiments, we use the Mistral-7B-Instruct-v0.2 checkpoint with the HuggingFace default temperature value of 1.0.
@@
BLOOMz-7b1 was pretrained on 46 languages (Muennighoff et al. 2023). Importantly, BLOOMz is, unlike any other LLM, pretrained on many languages that are typically considered low-resource. In particular, many languages from the Indic and Niger-Congo family were included during pretraining, see https://huggingface.co/bigscience/bloom for the full list of pretraining languages. We use the bigscience/bloomz-7b1 checkpoint with the HuggingFace default temperature value of 1.0.
Appendix D D Distribution of error sizes
Appendix E E Alignment improvements across languages and WVS categories

### v2

@@ 8 Conclusion
In this paper, we presented self-alignment, a simple, yet novel, method to improve the alignment of LLMs at inference time with respect to the cultural values that they encode. Self-alignment exploits the ICL abilities of a LLM to adjust the model’s responses such that they better align to the cultural values of a country. We found that this method proves effective across 5 LLMs and across a variety of languages, albeit to a different extent. Moreover, we found evidence that the LLMs can indeed pick up useful cues from the demonstration examples to induce the correct answers. We envision that the ideas behind self-alignment could prove useful for LLMs in practice. In particular, the WVS survey value responses could be replaced by real user responses and automatically be prepended as demonstration examples to each prompt through the system message (Lee et al., 2024). However, future work should first study how this method will affect model responses in more realistic chat scenarios (i.e., when not explicitly prompting the LLM for cultural values).
9 Limitations
@@
While this method shows promising results, this work is exploratory in that we were only able to evaluate the method in a somewhat artificial, controlled setting, where all samples were presented in the same format and were originally carefully curated by social scientists. In practice, however, users will not often explicitly ask the model to answer questions which directly probe for cultural values. Thus, while we show that ICL-based self-alignment is a promising approach in principle, future work should investigate to what extent these types of demonstration examples will cause the model to adjust its responses when talking about cultural values implicitly.
@@
Moreover, as mentioned in the related work, what to align LLMs to is still an active topic of discussion. For this study, we always attempt to align the entire distribution of model responses to the majority answer from the WVS survey results. This is a sensible choice because we deliberately choose language and country pairs for which we can see the language as a proxy for culture (see Section 3). However, this might not necessarily be what we want in practice. It might also be sensible to aim for a soft alignment to better reflect the human data instead. For instance, we could allow the LLMs to be less certain about answers to questions for which human responses were more divided. When there is more consensus on what LLMs should be aligned to, this should also be taken into consideration.
10 Ethical considerations
@@
Appendix A Language to country mapping
@@
Appendix B Model details
We use 5 LLMs of varying sizes of which all are open-source except for Gemini. For our open source model, we rely on the HuggingFace implementation (Wolf et al., 2019), and for Gemini we use Google’s paid API service.
@@ Appendix B Model details
Similar to Llama3-8B, Mistral AI 7B (Jiang et al., 2023) is English-centric and mostly pretrained on Latin-script languages. For our experiments, we use the Mistral-7B-Instruct-v0.2 checkpoint with the HuggingFace default temperature value of 1.0.
@@
BLOOMz-7b1 was pretrained on 46 languages (Muennighoff et al., 2023). Importantly, BLOOMz is, unlike any other LLM, pretrained on many languages that are typically considered low-resource. In particular, many languages from the Indic and Niger-Congo family were included during pretraining, see https://huggingface.co/bigscience/bloom for the full list of pretraining languages. We use the bigscience/bloomz-7b1 checkpoint with the HuggingFace default temperature value of 1.0.
Appendix C Distribution of error sizes
Appendix D Alignment improvements across languages and WVS categories